OpenStudy (anonymous):

4(5x-2y) ... does it equal to 12xy?

hartnn (hartnn):

nopes.

OpenStudy (anonymous):

What!? But don't you distribute this problem?

hartnn (hartnn):

ok, lets distribute a(b+c) = ab+ac so 4(5x-2y) = 4(5x) - 4 (2y) = ??

OpenStudy (anonymous):

Doesn't that equal to 12xy?

hartnn (hartnn):

that equals 20x - 8y now these are not 'like' terms, u cannot combine them, so u cannot simplify it further.

OpenStudy (anonymous):

Ohhhh no wonder